OpenStudy (anonymous):

how do you solve this equation using factoring x2+13x+40=0

OpenStudy (anonymous):

Solve using the quadratic equation: \[x=\frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\] x^2+13x+40 = 0 a=1, b=13, c=40

OpenStudy (anonymous):

Or by factoring: What are 2 factors (2 numbers that you multiply that will give you 40) of 40 (a*c=1*40=40) and when you add them together you get 13 (b). Can you find those factors?

OpenStudy (anonymous):

Are u there?

OpenStudy (anonymous):

yea thanks man im in algebra 2 class and these problems are killing me

OpenStudy (anonymous):

What are the factors of 40 2 numbers that you multiply that will give you 40? 1. 40 and 1 2. 2 and 20 Can you find the others?

OpenStudy (mew55):



OpenStudy (mew55):

x^2+8x+5x+40 x^2+13x+40